with airplane cruising 30,000ft, the navigator indicates to the captain that the plane should continue travelling north for 500 km and then turn to a heading of 45.0° east of north for 200 km. what will be the resultant distance traveled?
1
Expert's answer
2020-10-29T10:26:55-0400

The distance travel by the airplane in heading toward north

"d_1=500km"


The distance travel by the airplane in heading 45"\degree" east of north

"d_2=200km"


Resultant Distance travelled is given by


="\sqrt{d_1^2+d_2^2+2d_2d_2cos 45\degree}"


="\sqrt{500^2+200^2+2\times 500\times 200\times \dfrac{1}{\sqrt{2}}}"


="\sqrt{250000+40000+141400}"


="\sqrt{431400}=656.81km"


Hence The resultant distance is 656.81km